OpenStudy (anonymous):

What is the slope of the line parallel to -2y = 3x + 5

sam (.sam.):

-2y = 3x + 5 \[y = \frac{3x + 5}{-2}\] \[y = \frac{3x}{-2}+\frac{5}{-2}\]

sam (.sam.):

\[y = \frac{3x}{-2}+\frac{5}{-2}\] The coefficient of x is the gradient

OpenStudy (anonymous):

\[tilt \angle=\arctan(3/-2)\]

OpenStudy (anonymous):

-3

sam (.sam.):

How you get -3 @Kentie ?

OpenStudy (anonymous):

the slope of a parallel line is the opposite of the original slope.

OpenStudy (anonymous):

-1.5 is the answer or -2/3